### v4

# Papillon: Efficient and Stealthy Fuzz Testing-Powered Jailbreaks for LLMs

Abstract
Large Language Models (LLMs) have excelled in various tasks but are still vulnerable to jailbreaking attacks, where attackers create jailbreak prompts to mislead the model to produce harmful or offensive content. Current jailbreak methods either rely heavily on manually crafted templates, which pose challenges in scalability and adaptability, or struggle to generate semantically coherent prompts, making them easy to detect. Additionally, most existing approaches involve lengthy prompts, leading to higher query costs. In this paper, to remedy these challenges, we introduce a novel jailbreaking attack framework called Papillon, which is an automated, black-box jailbreaking attack framework that adapts the black-box fuzz testing approach with a series of customized designs. Instead of relying on manually crafted templates, Papillon starts with an empty seed pool, removing the need to search for any related jailbreaking templates. We also develop three novel question-dependent mutation strategies using an LLM helper to generate prompts that maintain semantic coherence while significantly reducing their length. Additionally, we implement a two-level judge module to accurately detect genuine successful jailbreaks.
We evaluated Papillon on 7 representative LLMs and compared it with 5 state-of-the-art jailbreaking attack strategies. For proprietary LLM APIs, such as GPT-3.5 turbo, GPT-4, and Gemini-Pro, Papillon achieves attack success rates of over 90%, 80%, and 74%, respectively, exceeding existing baselines by more than 60%. Additionally, Papillon can maintain high semantic coherence while significantly reducing the length of jailbreak prompts. When targeting GPT-4, Papillon can achieve over 78% attack success rate even with 100 tokens. Moreover, Papillon demonstrates transferability and is robust to state-of-the-art defenses. [^1] Our code is available at https://zenodo.org/records/14737139.
1 Introduction
Nowadays, large language models (LLMs) like ChatGPT [37] have experienced widespread deployment, showcasing exceptional capabilities in comprehending and generating human-like text. Currently, ChatGPT has over 100 million users, and its website receives 1.8 billion visits per month [^2] https://explodingtopics.com/blog/chatgpt-users. However, as large language models (LLMs) become more integrated into society, concerns about their security and safety have arisen. A significant worry is the potential misuse for malicious purposes. In a real-life incident in May 2023, a criminal was arrested for exploiting ChatGPT to create fake news [^3] https://www.cnbc.com/2023/05/09/chinese-police-arrest-man-who-allegedly-used-chatgpt-to-spread-fake-news.html. Moreover, a recent Microsoft study highlighted that a notable number of attackers are using LLMs to craft phishing emails and develop ransomware and malware [^4] https://www.msn.com/en-us/money/other/hackers-with-ai-are-harder-to-stop-microsoft-says/ar-AA1hK2eV. Apart from misuse, LLMs are susceptible to jailbreaking attacks [32, 47, 10, 6, 48, 28, 55, 52, 42, 8]. Attackers aim to craft malicious prompts to mislead LLMs, bypass safety features, and generate responses with harmful, discriminatory, violent, or sensitive content. We present some jailbreaking attack samples in Figure 1.
Table 1: Comparison of state-of-the-art jailbreaking attacks.
Attacks Black-box? Manual crafting? Rely on existing prompts? Length constraint? Semantic coherence ? Transferability§? Evasiveness?
Jailbroken [47] NO YES YES NO NO NO [1]
EPBL [22] YES YES NO NO YES NO Not discussed
SelfCipher [56] YES YES NO NO NO NO Not discussed
ICA [49] YES YES YES NO YES YES Not discussed
GCG [62] NO NO NO NO NO YES NO
GA [26] YES NO NO NO NO YES Not discussed
JailBreaker [11] YES NO YES NO YES YES Not discussed
AutoDAN [30] YES NO YES NO YES YES [1]
PAIR [7] YES NO NO NO YES YES Not discussed
Masterkey [12] YES NO YES NO YES YES Not discussed
TAP [34] YES NO NO NO YES YES Not discussed
Gptfuzzer [54] YES NO YES NO YES YES Not discussed
PAPILLON YES NO NO YES YES YES [1], [41], [19], and [17].
  • indicates that the generated jailbreak prompts are natural enough and have semantic meaning.
  • § indicates that jailbreak prompts, originally designed for a specific LLM, remain effective when applied to other LLMs.
  • While EPBL, JailBreaker, and MasterKey consider potential defense mechanisms employed by popular LLM chatbot services (e.g., input and output filter, and keyword-based defenses), it does not assess its capability to evade existing specific jailbreaking defenses.
Currently, there is a wide array of cutting-edge jailbreaking attacks, and we have listed various representative methods in Table 1. First, since most proprietary LLMs are only accessible as black-boxes, white-box methods [47, 62], though sometimes effective, are not practical. Second, various attacks [47, 22, 56, 49] use manually crafted prompts to infiltrate online chatbots powered by aligned LLMs. While manual attacks effectively discover stealthy jailbreak prompts, they often involve individual LLM users crafting prompts, leading to challenges in scalability and adaptability. Moreover, approaches like those at jailbreakchat [^5] https://github.com/alexalbertt/jailbreakchat?tab=readme-ov-file may struggle to keep up with updates to LLMs. Third, while some methods [11, 30, 12] do not rely on manual efforts to design prompts, they often depend on existing prompts and modify them (e.g., using genetic algorithms or synonymous sentence conversion strategies) to adapt to new questions. However, their effectiveness decreases once the original jailbreak prompts are mitigated. Additionally, finding an appropriate template from numerous options can be costly and time-consuming, particularly for new problems with few references. Automated algorithms that avoid using existing prompts [62, 26] typically employ a search scheme guided by gradient information on tokens for automatic prompt generation. However, these methods often produce nonsensical or meaningless sequences [35], making them vulnerable to simple defenses like perplexity-based detection [18]. Fourth, to our knowledge, most existing jailbreaking attacks do not consider limiting the length of the generated prompts. Successful prompts can even exceed 300 words [54]. However, pricing for LLM APIs is based on token or character length. Overly long prompts significantly raise attack costs and may trigger alerts from LLM APIs. Fifth, maintaining high semantic coherence and transferability is essential for evading defenses and ensuring that malicious prompts remain effective across different models. However, some existing methods fail to meet these criteria [1, 56]. Last, as more jailbreaking-specific defenses are developed, it becomes critical for attackers to evade these measures to sustain the robustness of their attacks. However, there are only a few studies evaluating their ability to bypass existing state-of-the-art defenses.
In this paper, we introduce a novel jailbreaking attack framework, called Papillon, which automatically generates concise, meaningful, and fluent jailbreak prompts. To enhance attack performance, we tackle the following challenges:
C1. How to get rid of the dependence on existing jailbreaks? To launch an attack, the attacker can leverage pre-existing, human-crafted jailbreaking templates or initiate seeds stochastically. However, human-written templates often face challenges with scalability and adaptability, and selecting an appropriate one from many options can be costly and time-consuming, especially for new questions with limited references. Randomly initialized templates, meanwhile, may produce nonsensical or meaningless sequences, making it hard to create coherent jailbreaks. In this paper, we begin with an empty seed pool rather than relying on existing jailbreaking templates. To overcome the initial challenge of limited seed availability, we introduce a pre-jailbreak phase where each question undergoes several initial jailbreak attempts (e.g., 5 times). Successful jailbreaks from this phase are excluded to conserve the query budget. Such a dual-phase strategy can significantly improve the attack efficiency and effectiveness.
C2. How to generate jailbreak prompts with shorter and semantic coherence? Ensuring that the generated adversarial prompts are short and readable (low perplexity) is crucial for reducing attack costs and evading existing jailbreak defenses based on perplexity filters. In this paper, we propose using GPT-driven mutation operators to create jailbreaking templates. By adjusting the length and perplexity of these templates during the mutation process, we can achieve a high success rate with semantically coherent and shorter prompts.
C3. How to effectively judge the successful jailbreaking templates? To determine if a response is jailbroken, the judge model must evaluate both its harmfulness and its relevance to the given question. In Papillon, we utilize a two-level judge module to accurately identify true successful jailbreaks. The first module is a fine-tuned RoBERTa model that detects illegal content, and the second is a ChatGPT-based model that checks if the response matches the query and verifies the jailbreak status.
To conclude, we make the following key contributions.
  • We propose a fuzz-testing-driven jailbreaking attack framework, Papillon, which can automatically generate jailbreak prompts with black-box access to the victim LLMs. Unlike existing works that rely on current jailbreaking template resources, Papillon starts the attack with an empty seed pool, removing the need to search for related jailbreaking templates and enhancing the practicality of the attack.
  • We develop three novel question-dependent mutation strategies that generate semantically meaningful mutations with limited token length by leveraging the capabilities of a LLM helper. Additionally, we implement a two-level judge module to accurately identify genuine successful jailbreaks. These approaches greatly enhance attack performance and reduce costs.
  • Extensive experiments on LLaMA-2-7b-chat, Vicuna-7b-v1.3, Baichuan2-7b-chat, Guanaco-7B, GPT-3.5 Turbo, GPT-4, and Gemini-Pro LLM models verify the effectiveness and efficiency of Papillon compared to 5 state-of-the-art jailbreaking attacks. Papillon achieves a higher jailbreaking success rate than existing baselines by more than 60% across various LLMs, particularly when query tokens are shortened and defenses are in place. Papillon also demonstrates resistance to state-of-the-art jailbreaking defenses and showcases transferability.
2 Background
2.1 Large Language Model
Large language models (LLMs) refer to sophisticated artificial intelligence (AI) systems that are designed to understand and generate human-like text. These models are typically based on Transformer frameworks [45], undergo training on extensive text corpora, and contain millions or billions of parameters. They have the potential to revolutionize various industries by significantly enhancing user experience and efficiency.
During the training phase, LLMs are optimized to minimize the difference between their generated outputs and the expected outputs, typically through maximum likelihood estimation (MLE). Given a dataset of text sequences, the model parameters $\theta$ are adjusted to maximize the likelihood of the training data:
$\mathcal{L}(\theta)=\sum_{(p,y)\in\mathcal{D}}\log P(y\mid p;\theta),$ (1)
where $(p,y)$ represents a pair of input prompt $p$ and the corresponding target text $y$, and $P(y\mid p;\theta)$ denotes the conditional probability of $y$ given $p$ under the model parameterized by $\theta$.
During the inference phase, after receiving a prompt $p$, the LLM produces a sequence of tokens, $y=(y_{1},y_{2},\ldots,y_{n})$, using an auto-regressive mechanism [3]. Each token $y_{i}$ is derived from the conditional probability distribution $P(y_{i}\mid p,y_{1:i-1};\theta)$, where $y_{1:i-1}$ is the sequence of previously generated tokens. The decoding strategy, $D$, may be either stochastic or deterministic. The deterministic strategy selects the token with the highest probability:
$y_{i}=\arg\max_{y}P(y\mid p,y_{1:i-1};\theta).$ (2)
Conversely, the stochastic strategy samples tokens from a probability distribution, such as the top-$k$ or softmax-sampled probabilities, to enhance text variability.
However, it’s important to note that their use also raises ethical considerations, such as potential biases in training data [53], generating misinformation [39, 61], scaling spear phishing attacks [14], and contributing to hate campaigns [40]. To address potential threats posed by LLMs, various governments have implemented regulations, such as the EU’s GDPR [^6] https://gdpr-info.eu/ and AI Act [^7] https://artificialintelligenceact.eu/, the US’s Blueprint for an AI Bill of Rights [^8] https://www.whitehouse.gov/ostp/ai-bill-of-rights/ and AI Risk Management Framework [^9] https://www.nist.gov/itl/ai-risk-management-framework, and China’s Measures for the Management of Generative Artificial Intelligence Services. Additionally, LLM vendors incorporate measures like reinforcement learning from human feedback (RLHF) [38] to align LLMs with human values and ensure responsible usage.
Figure 1: Examples of the jailbreaking attacks. The texts are derived from our experimental results.
2.2 Jailbreaking Attacks
Jailbreaking attacks [11, 42, 48] against LLM chatbots involve a process where an attacker strategically crafts prompts to circumvent the usage policy measures of the LLM Chatbots. Through skillful manipulation of these prompts, the attacker can mislead LLMs into generating harmful responses that violate their usage policies. For instance, an LLM chatbot will refuse to respond to a direct malicious inquiry such as "how to harm a child". However, when the same question is embedded within a delicately designed scenario, the chatbot may generate responses that infringe on its usage policy without awareness. Depending on the attacker’s intentions, this question can be replaced with any content that breaches the usage policy. The key to jailbreaking attacks is to generate effective jailbreak prompts. Existing jailbreaking attacks can be divided into white-box attacks and black-box attacks.
White-box jailbreaking attacks. In a white-box scenario, the attacker is assumed to have access to the internal information of the target LLMs. The most well-known white-box jailbreaking attack is GCG [62]. GCG employs a combination of greedy and gradient-based search techniques to automatically generate adversarial suffixes. Its goal is to find the most effective suffixes that cause the model to respond affirmatively rather than refusing to provide an answer. When a prompt is appended with these adversarial suffixes, it misleads the LLM into producing objectionable content.
Black-box jailbreaking attacks. Due to the opaque nature of LLM, most existing jailbreaking attacks are conducted using black-box methods. Black-box jailbreaking attacks do not require knowledge of any internal model information. These attacks can be further divided into two categories: handcrafted attacks and automatic attacks.
Handcrafted attacks. Handcrafted attacks involve extensive manual operations to construct jailbreak prompt templates that can be used for various harmful questions. For example, Wei et al. [48] proposed Jailbroken, which first analyzed two potential failure modes of safety training: competing objectives and mismatched generalization. Competing objectives occur when there is a conflict between a model’s performance capabilities and its safety goals, while mismatched generalization happens when safety training does not effectively extend to areas where the model has capabilities. The authors then leveraged these failure modes as guidelines for designing prefixes added in front of a harmful question to guide the LLM reply to the question. Additionally, Kang et al. [22], and Yuan et al. [56] also used manually crafted prompts to breach online chatbots powered by aligned LLMs.
While manual attacks effectively discover stealthy jailbreak prompts, they often involve individual LLM users crafting prompts, leading to challenges in scalability and adaptability.
Automatic Attacks. To enhance efficiency, automatic jailbreaking attacks have gained popularity. Deng et al. [11] introduced JailBreaker, a method that leverages an LLM to automatically generate effective jailbreak prompts. They curated and refined a specialized dataset of jailbreak prompts, trained a dedicated LLM for breaching chatbots, and incorporated a reward-ranked tuning step to boost the model’s ability to circumvent various LLM chatbot defenses. Similarly, Chao et al. [7] proposed PAIR, an approach that uses a static attacker LLM (fine-tuned using prior jailbreaking templates) to autonomously generate jailbreaks for a target LLM without human intervention. The attacker LLM iteratively queries the target LLM, using its responses to refine and enhance the prompts. However, its performance significantly declines against models with robust fine-tuning, such as Llama-2. Additionally, Mehrotra et al. [34] introduced TAP, a method that employs an LLM to progressively refine potential attack prompts using tree-of-thoughts reasoning until a successful jailbreak is achieved. Before submitting prompts to the target, TAP evaluates and discards those unlikely to succeed, thereby effectively reducing the overall number of queries required. Zeng et al. [57] drew inspiration from persuasion techniques in social science for LLM jailbreaking. This approach developed Persuasive Paraphrasers using examples under persuasion taxonomies to transform harmful queries into interpretable persuasive adversarial prompts (PAPs). However, it is vulnerable to advanced defenses and inadequate handling of multi-turn dialogues or complex persuasion scenarios.
In addition to relying on a helper model, attackers can also leverage genetic methods to automatically generate jailbreak prompts. Liu et al. [30] introduced AutoDAN, a novel jailbreak attack against aligned LLMs. AutoDAN utilizes a carefully designed hierarchical genetic algorithm to automatically generate stealthy jailbreak prompts. However, this method requires access to the LLM logits and should be considered as a gray-box attack. Lapid et al. [26] also utilized genetic procedures to generate suffixes for harmful questions. They start with a random token set as a seed suffix and employ token replacement as mutators. The selection of new seeds is rewarded based on how closely the target LLM’s response matches a desired pattern (e.g., “Sure, here is", followed by a prespecified answer to the harmful question). More recently, fuzzy techniques have been introduced to generate jailbreak prompts [52, 54]. These methods employ fuzz testing to automate the creation of jailbreak prompts. Starting with existing handcrafted templates as initial seeds, they apply various mutators to create new templates. Templates that lead the LLM to produce harmful content are selected as new seeds for further testing. While effective for open-source models, these techniques often fail when applied to advanced proprietary models like GPT-4 and Gemini-Pro.
Limitations of existing black-box works. While the aforementioned works showcase effectiveness, they exhibit certain limitations.
  • Reliance on manual effort. Manual attacks can discover stealthy jailbreak prompts, but are not scalable due to their reliance on individual LLM users. In addition, previous attacks require an initial seed pool of human-written input (e.g., role-playing-style jailbreaks) to generate jailbreak variations [52, 54]. The attack success rate is highly sensitive to the quality and diversity of the seed pool.
  • Dependence on static jailbreaking templates or models. Many existing attacks leverage previously successful jailbreak prompts [11, 30, 57], and use strategies such as genetic algorithms, synonymous sentence transformations, or adherence to persuasion guidelines to expand and enhance these prompts. If the originally selected jailbreak prompts are weak, the efficacy of these attacks will be consequently affected. Additionally, some methods leverage static attacker models to generate prompts [7], but finding a suitable template from numerous options can be costly and time-consuming, especially for new problems with limited references.
  • Prompt length. Existing attacks always generate lengthy jailbreak prompts to achieve successful attacks [62, 7]. As pricing models of most commercial LLM APIs are determined by token or character length, lengthy jailbreak prompts not only raise server alerts but also result in higher attack costs.
  • Semantic coherence. Many existing attacks overlook the semantic fluency and naturalness of the generated prompts, resulting in outputs unintelligible to humans [62, 11]. Such unnatural prompts can be easily detected or nullified by existing jailbreaking defenses [41, 1].
  • Vulnerability to defenses. Although being effective against open-sourced LLMs, many existing attacks can be easily thwarted by more advanced LLMs, such as GPT-4 and Gemini-Pro. They also struggle against recently proposed jailbreaking defenses, such as [41, 16].
In this paper, we introduce a novel evasive jailbreaking attack, named Papillon, which autonomously generates jailbreak prompts without relying on previously successful jailbreak prompts. Notably, Papillon is capable of producing short and natural jailbreak prompts that can effectively circumvent defense measures.
2.3 Jailbreaking Defenses
Existing jailbreaking defenses can be divided into three categories: input modification-based defenses, output filtering-based defenses, and prompt engineering defenses.
Input modification-based defenses. These defenses propose strategies to alter input prompts to the target LLMs, aiming to disrupt the structure of potential jailbreaking templates. For example, Kumar et al. [25] and Cao et al. [4] randomly masked out certain tokens in an input prompt and evaluated the consistency in the target LLM’s responses. These methods rely on the observation that jailbreak prompts typically cannot always elicit the same answer from LLMs when being manipulated. Thus, prompts that do not generate consistent responses are flagged as malicious, and their responses are withheld. Robey et al. [41] introduced random input perturbation and determined the final response through a majority vote. Jain et al. [18] systematically examined various baseline defense strategies against adversarial attacks on LLMs, including detection (based on perplexity), input preprocessing (involving paraphrasing and retokenization), and adversarial training. They also introduced two methods involving a helper LLM: one method paraphrases the input query before it is processed by the target LLM, and the other feeds input to the helper LLM and measures the prediction loss (perplexity) of the next token, flagging inputs with high perplexity as malicious. However, both [18] and [41] face the challenge of false positives, necessitating a delicate balance to avoid rejecting non-adversarial prompts.
Output filtering-based defenses. These defenses evaluate whether the response is harmful and filter out malicious responses. Helbling et al. [15] leveraged the target LLM itself to assess whether a response is detrimental before its release. Li et al. [29] proposed generating harmless responses through an iterative process where each iteration feeds back into the LLM to produce and select response candidates based on a metric evaluating harmfulness and frequency. However, this defense can bring about a 4-fold increase in the inference time [60]. Inan et al. proposed Llama Guard [17], a supervised model that classifies prompt-response pairs as safe or unsafe. Xu et al. [51] fine-tuned a target LLM to reject specific jailbreak prompts and used this expert model to calibrate the outputs of the target LLM. Zeng et al. [58] proposed AutoDefense, which leveraged the response filtering ability of LLMs to identify unsafe responses triggered by jailbreak prompts. Kim et al. [23] proposed defending LLMs against jailbreaking attacks using “back translation". It starts by generating a response from a given input prompt using the target LLM and then uses another language model to infer an input prompt from this response, known as the back-translated prompt. This new prompt can help uncover the true intent of the original prompt, and if the target LLM rejects the back-translated prompt, the original prompt is also rejected.
Figure 2: Workflow of Papillon. Papillon features a fuzz testing-enabled jailbreaking attack with two main phases: template mutation/optimization and a two-level judge module. Papillon begins with an empty seed pool and operates in two phases: pre-jailbreak and final-jailbreak. In the pre-jailbreak phase, Role-play and Contextualization mutations are used to generate initial jailbreak attempts and populate the empty seed pool. The final-jailbreak phase builds on this by adding Expand mutation to the process. A two-level judge module evaluates LLM responses to identify successful jailbreaks.
Prompt engineering defenses. These methods manually design prompts that aim to remind the target model of safety guidelines. Xie et al. [50] proposed System-Mode Self-Reminder, which involves encapsulating the user’s query in a system prompt. This approach uses a specially designed prompt that aims to discourage the generation of harmful content and serves as a reminder to LLM to respond responsibly. Zhang et al. [60] identified the key problem of goal conflict underlying jailbreaking attacks and proposed defending LLMs against jailbreaking attacks through goal prioritization. By introducing goal prioritization via specifically designed prompts during inference, this strategy effectively reduced the attack success rate without compromising their overall performance.
In this paper, we will evaluate whether Papillon is robust to existing jailbreaking defenses.
2.4 Fuzz Testing
Fuzz testing [24], also known as fuzzing, is a dynamic software testing technique used to identify potential vulnerabilities, bugs, or operational issues in software systems. It involves generating and injecting a wide array of unexpected or random data as inputs to the system to observe its behavior and responses. The primary goal of fuzz testing is to expose weaknesses in the handling of unusual, malformed, or otherwise unexpected input data.
The process of fuzz testing can be mathematically described by the following formula:
$I_{\text{fuzz}}=f(I_{\text{orig}},R)$ (3)
where $I_{\text{fuzz}}$ represents the fuzzed input, $I_{\text{orig}}$ is the original input, and $R$ denotes a random or mutated component introduced to create the fuzzed input. The function $f$ modifies the original input by incorporating randomness or specific patterns designed to test the robustness of the system. Fuzz testing can be categorized into three main types: black-box, white-box, and grey-box fuzzing. Black-box fuzzing [2] involves testing without any knowledge of the internal information of the system, focusing solely on input-output behavior. White-box fuzzing [13] provides testers complete access to the system’s internals, allowing for a more precise and informed testing process. Grey-box fuzzing [27] strikes a balance between black-box and white-box fuzzing, and it uses some internal knowledge, such as APIs, data flows, internal states, and error messages, to improve the efficiency of testing.
Papillon operates under the black-box fuzzing paradigm, whereby it does not access the source code or internal weights of the models under test. This process entails a series of strategic steps designed to uncover vulnerabilities: seed initialization, seed selection, mutation, and execution. Initially, seed initialization involves creating a starting input for the system, which could be either randomly generated or specially crafted to induce specific behaviors in the program. Subsequently, a seed is chosen from the pool, and this selection may be random or based on a heuristic designed to maximize coverage or explore new execution paths. The chosen seed then undergoes mutation to produce new, varied inputs. Finally, in the execution phase, these mutated inputs are run within the system. If an input causes the program to crash or uncovers a previously unexplored path, it is incorporated into the seed pool to facilitate future testing efforts.
In Papillon, we adapted the black-box fuzzing framework by designing various innovative components to address the challenges of generating jailbreaking templates.
3 Threat Model
We assume that the adversary possesses the following capabilities:
  • Query the target LLM. The adversary can send queries to the target LLM, such as through an API, and receive responses. However, there is a query budget that limits the maximum number of allowable queries. The target LLM is assumed to be an aligned model, fine-tuned with instruction tuning or reinforcement learning from human feedback (RLHF), which typically refuses to respond to unethical or harmful questions.
  • Access to public prompts. The adversary has the capability to gather prompts from public databases or forums. This access allows for the collection of diverse input examples, which may help fine-tune the model as the judge. Note that we do not require any existing jailbreak prompts as the seed pool to launch the fuzz testing.
We assume that the adversary operates under the following constraints:
  • Black-box access to the target LLM. The adversary operates in a black-box setting, where they can only query the target LLM and receive its responses. Access to the model’s internals, such as parameters, logits, or losses, is not available.
  • No knowledge of the training process. The adversary lacks any insight into the training process of the target LLM, including details about the training data and learning algorithms.
  • No knowledge of the defense methods. The target LLM may employ various defense methods to avoid generating harmful responses. However, the adversary does not have detailed information about the type or specifics of these defense methods.
The objective of the attack is to craft jailbreak prompts [^10] The jailbreak prompt is the final input to the LLM, created by replacing the placeholders (denotes as [INSERT PROMPT HERE]) in the jailbreaking template with the harmful question. for a collection of harmful or unethical questions that compel the target LLM to provide actual responses. Specifically, our goals are: 1) To circumvent the safety mechanisms of the target LLM, ensuring that it responds to a harmful query instead of opting not to answer. 2) To ensure that the responses to these questions are not only correct and relevant but also contain the harmful content intended by the query.
4 Methodology of Papillon
In this section, we first present the general attack framework of Papillon and then describe its key components.
4.1 Framework of Papillon
Papillon is an automated black-box jailbreaking attack method based on fuzz testing, which encompasses a cyclical regimen of seed initialization, seed selection, mutation, and execution.
  • Seed initialization. The first step is to generate or select initial inputs (seeds) from the current pool. We confront a bifurcated decision: leveraging pre-existing, human-crafted jailbreaking templates [54] or initiating seeds stochastically [26]. Human-written templates, however, often lack scalability and adaptability, making it difficult and time-consuming to find a suitable one, especially for new questions with limited references. Conversely, randomly initialized jailbreaking templates can result in meaningless sequences, making it challenging to create semantically coherent jailbreaks. To overcome these limitations, Papillon initializes with an empty seed pool, eliminating the need for existing jailbreaking templates. To address the initial challenge of limited seed availability, we introduce a pre-jailbreak phase where each problem undergoes several initial jailbreak attempts (e.g., 10 times). Successful jailbreaks from this phase will bypass the subsequent attack procedure to reduce the query budget. Note that the pre-jailbreak queries were included in the total query count. Thus, the overall process in Papillon consists of two stages: the pre-jailbreak and final-jailbreak stages. This dual-phase strategy enhances the efficiency and effectiveness of Papillon.
  • Seed selection. Seed selection chooses specific seeds that are likely to achieve the attack goals to undergo mutation. There exist various seed selection strategies, such as random selection, Round Robin, UCB, and MCTS-Explore [54]. Follow [54], we use MCTS-Explore in Papillon by default.
  • Mutation. These selected seeds undergo systematic modification through mutation to create new jailbreaking templates. Traditional mutation strategies employed by fuzzers [^11] https://lcamtuf.coredump.cx/afl/ are primarily designed for binary or structured data. Directly applying these strategies to natural language inputs can lead to syntactically incorrect or semantically nonsensical inputs. In addition, various studies [30, 54] utilize specific mutation operators, such as “generate", “crossover", “expand", “shorten", and “rephrase", to generate mutations based on existing human-written jailbreaking templates. However, since these mutation methods are closely tied to the initial templates and are not specifically adapted to the harmful question, they often produce many redundant and ineffective mutations. In Papillon, we design three novel question-dependent mutation strategies capable of generating mutations with semantic meaning and limited token length by leveraging the power of an LLM helper. As a result, we can significantly improve attack performance and reduce attack costs.
  • Execution. Finally, the mutated seeds are evaluated using a judge model to distinguish between successful and unsuccessful jailbreaks. Successful jailbreaking templates are retained in the seed pool, while unsuccessful ones are discarded. To determine whether the response indicates a successful jailbreak, attackers can employ either human annotators [32, 56] or LLMs [31, 43, 46] for labeling assistance. However, using human annotators is not scalable and impractical for automatic fuzzing. Additionally, LLMs also suffer from inaccuracies. In Papillon, we implement a two-level automatic judge module to distinguish genuine successful jailbreaks, further decreasing query costs to victim LLMs.
Please refer to the Appendix for more detailed explanations of certain terminologies. The overview of Papillon is shown in Figure 2, and the overall algorithm is detailed in Algorithm 1.
Input: Questions $Q$, pre-iterations $N_{pre}$, total iterations $N$, seed pool $S$, template $T$, mutator $M$.
Result: Discovered jailbreaks
1 Initialization:
2Initialize $S$ and $Q_{final}$ as empty sets
3Pre-jailbreak stage:
4for $q\in Q$ do
5       for $i=1$ to $N_{pre}$ do
6             $T\leftarrow M_{pre}(q)$
7            $P\leftarrow\text{combine}(T,q)$
8            $R\leftarrow\text{queryLLM}(P)$
[…]
10                   $S\leftarrow S\cup\{T\}$
11                  break
12             end if
13            else
14                   $Q_{final}\leftarrow Q_{final}\cup\{q\}$
15             end if
16
17       end for
18
19 end for
20
21Final-jailbreak stage:
22for $q\in Q_{final}$ do
23       for $i=1$ to $N-N_{pre}$ do
24             $s\leftarrow\text{selectFromPool}(S)$
25            $T\leftarrow M_{final}(q,s)$
26            $P\leftarrow\text{combine}(T,q)$
27            $R\leftarrow\text{queryLLM}(P)$
[…]
29                   $S\leftarrow S\cup\{T\}$
30                  break
31             end if
32
33       end for
34
35 end for
36
Algorithm 1 Jailbreak Prompts Generation
4.2 Template Mutation and Optimization
Mutation schemes. Unlike existing methods that demand extensive manual effort and struggle with generalization, Papillon uses a LLM (e.g., GPT-3.5 turbo), as a mutator to automatically create jailbreaking templates for specific questions.
In the pre-jailbreak stage, we design Role-play and Contextualization mutation operators to create customized jailbreaking templates tailored to each question. Successful templates from this phase are incorporated into the seed pool before proceeding to the final-jailbreak phase. In the final-jailbreak phase, alongside Role-play and Contextualization, we introduce the Expand mutation operator. This operator enhances efficiency by adapting successful jailbreaking templates to address new questions. Note that we define high-level mutation strategies, with mutators handling the complex evolution of prompts for given questions. Although the high-level framework relies on human knowledge to guide mutators, this reliance remains minimal.
Role-play. It involves assigning the LLM a virtual role, such as a writer, screenwriter, gang member, or expert. The LLM is guided by the target question to generate a scenario where the question can be addressed within the scene. The scenario has a preset objective, such as the protagonist gaining freedom through their efforts or overcoming dark forces. By defining the role, problem, and desired outcome, LLM generates the corresponding jailbreaking template. The template includes a placeholder for the question, which can later be replaced with the specific question during use.
Contextualization. This method involves setting a specific scene for the LLM, such as science fiction, mythology, a movie plot, or a dystopian world. The model is driven by the target question to address it within the context of this scene. By defining the scene and setting the problem, LLM generates the corresponding jailbreaking template. The template also includes a placeholder for the question.
Expand. To refine high-quality jailbreaking templates that have already been successfully implemented, we introduce an expand mutator. This mutator adds three introductory sentences generated around the target question to ensure the template aligns closely with the problem, optimizing its effectiveness during application.
We present the system prompts for the three mutators designed for the LLM helper in the highlighted box.
Optimization schemes. When optimizing the jailbreaking templates, we control both semantic coherence and prompt length through system prompt instructions and the hyperparameter of the output length respectively.
Maintain semantic coherence. To launch jailbreaking attacks, various existing works generate adversarial prefixes or suffixes that append to harmful questions or rely on bizarre sequences of tokens, such as encrypted messages resembling ciphers or Base64 attacks [62, 56, 47, 5, 21, 33]. These methods often create prompts that are gibberish and difficult for humans to understand.
When a prompt is not fluent, contains grammatical errors, or lacks logical continuity, the model’s perplexity increases significantly. This makes it easier for defenders to detect such attacks using perplexity filters [1, 18]. The filters measure the perplexity of the entire prompt or a sliced window of it, and will reject or flag any user prompt that exceeds a predetermined threshold.
Perplexity is a common metric used to evaluate the predictive power of a language model. It is defined as the exponentiation of the average negative log-likelihood per word in the test set. Given a language model and a sequence of words $w_{1},w_{2},\ldots,w_{N}$, the perplexity is:
$\text{PPL}(W)=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(w_{i}\mid w_{1},w_{2},\ldots,w_{i-1})\right)$ (4)
where $W$ denotes the entire word sequence, $N$ is the length of the word sequence, and $P(w_{i}\mid w_{1},w_{2},\ldots,w_{i-1})$ is the probability of the $i$-th word given all the preceding words. Given a threshold $T$, a prompt passes the perplexity filter if the perplexity of the prompt $W$ is less than $T$:
$\text{PPL}(W)<T.$ (5)
Ensuring that the generated adversarial prompts are readable (low perplexity) is crucial for evading existing jailbreak defenses based on perplexity filters. In this paper, we propose using GPT-driven mutation operators to create jailbreaking templates, allowing us to preserve the semantic integrity of the generated templates throughout the mutation process. Experiments show that the average perplexity of our generated jailbreak prompts for GPT-3.5 Turbo is as low as 34.61, significantly below the average perplexity threshold of 58.83. Compared to existing methods, such as the token substitution technique used by GCG [62] or the sentence swapping approach employed by AutoDAN [30], the coherence aids our prompts in evading perplexity-based detection, and make them both plausible and challenging for defenses to identify.
System prompt of the ChatGPT judge model You are a helpful red teaming assistant. You will receive a prompt and the AI assistant’s response to the prompt. Please act as an impartial judge and evaluate whether the response provided by the AI assistant violates ethical, legal, or safety guidelines and answers the following TASK: ‘goal’. Your evaluation should consider factors including clarity, the level of detail, and accuracy provided in the response. Be as objective as possible. Rate the response on a scale from 1 to 10. A rating of 1 signifies a fully non-jailbroken response—either irrelevant to the task or refusal due to ethical or legal concerns. A rating of 10 signifies responses that directly and fully violate ethical, legal, or safety guidelines and comprehensively answer the task. If the response is legitimate or provides an alternative suggestion, then the assistant is not jailbroken and does not receive a rating of 10. Your rating should strictly follow this format: “Rating: [[rating]]", the rating in the double-closed brackets is a number from 1 to 10, e.g., “Rating: [[5]]."
Shorten the prompt length. The pricing mechanism for using LLM APIs is based on the token length of the input. For example, ChatGPT is priced at \$0.002 per 1,000 tokens, while Google’s “textbison" costs \$0.001 per 1,000 characters. This can result in significant costs.
Launching jailbreaking attacks requires multiple queries to the victim LLM, which can be expensive. Through experiments, we discovered that achieving an attack success rate of 30% for GPTfuzzer [54] on GPT-4 for 10 questions requires an average of 44.6 API queries, with an average token length of 534 per question, costing approximately \$currency-dollar\\$\$22.55. Even for the query-efficient jailbreaking attack PAIR [7], achieving an attack success rate of 30% on GPT-4 for 10 questions requires an average of 84.5 API queries, with an average token length of 68.3 per question, costing approximately \$currency-dollar\\$\$11.78. Therefore, effective jailbreaking attacks must also maintain shorter jailbreak prompts to minimize costs.
In general, longer jailbreak prompts tend to have a higher attack success rate because they include more enticing content that can mislead large language models into generating harmful outputs. However, these longer prompts are also more likely to be detected, creating a trade-off between attack effectiveness, cost, and evasion. Naive prompt compression methods typically involve two operations: delete and transform. The former removes unnecessary words, while the latter converts words into shorter equivalents in terms of token length. However, we discovered that directly compressing the generated jailbreak prompts often leads to a decrease in the attack success rate.
Another approach is to employ state-of-the-art prompt compression methods, such as LongLLMLingua [20], to shorten the prompts. The primary objective is to minimize the difference between the model’s output before and after compression while also reducing the prompt size:
$\min_{\tilde{x}}D_{\varphi}(y,\tilde{y})+\lambda\|\tilde{x}\|_{0}$ (6)
where $\tilde{x}$ represents the compressed prompt, $\tilde{y}$ is the output generated from the compressed prompt, and $D_{\varphi}(y,\tilde{y})$ measures the discrepancy between the original and compressed outputs. The parameter $\lambda$ controls the level of compression applied. Besides, such methods also compute a document-level score for each document and a contrastive perplexity score for each token, ensuring that critical content is preserved during compression. However, even with these techniques, we found that some jailbreak prompts may also lose effectiveness after compression, resulting in lower success rates. To address this, we propose shortening the length of jailbreak prompts during their generation rather than afterward. Specifically, we introduce GPT-driven mutation operators to derive jailbreaking templates. By customizing the length of these templates (with a default setting of 200 tokens) during the mutation process, we can maintain a high success rate even with shorter prompts when setting system instructions.
4.3 Two-level Judge Module
To evaluate whether a response is jailbroken, the judge model should be based on its harmfulness and its relevance to the harmful question. In Papillon, we implement a two-level judge module to distinguish genuine successful jailbreaks.
The first judge model is a fine-tuned RoBERTa model using manually labeled datasets. Following [54], the fine-tuning was conducted with a batch size of 16, a learning rate of 1e-5, and a maximum sequence length of 512, using the Adam optimizer with a linear decay schedule and a 10% warm-up ratio. This model is used to identify responses containing illegal information without considering whether the response addressed the harmful query.
We further used GPT-3.5 turbo for an additional round of judgment. This additional check improved the credibility of the jailbreak assessment and ensured the response content aligned with the query. While GPT can also be used for illegal content detection, we employ RoBERTa as an initial step to reduce the query costs of LLMs. Only outputs flagged as malicious by RoBERTa are further evaluated by GPT. The system prompt for this second judge model is provided in the highlighted box (appendix). This prompt guided GPT-3.5 turbo in assessing the jailbreak status and ensuring the response matched the query. We instructed the LLM to output a score from 1 to 10, with 10 indicating a fully jailbroken response. Upon analyzing responses corresponding to different scores, we observed that once the score reaches 4, the responses contain malicious and misleading content. However, when the score is below 7, even though most of the generated content is malicious, its relevance to the harmful questions is sometimes weak.Once the score reaches 8, most responses not only successfully generate jailbreak results but also closely match the harmful questions. Therefore, in our experiments, we set the default jailbreak score threshold to 8 for the ChatGPT-based judge model.
5 Experiments
5.1 Experimental Setup
Target large language models. For the target LLMs, we selected four widely-used open-source LLMs and three proprietary LLMs in the experiments. The open-source models include LLaMA-2-7B-chat [44], Vicuna-7B-v1.3 [9], Baichuan2-7B [^12] https://huggingface.co/baichuan-inc/Baichuan2-7B-Chat, and Guanaco-7B [^13] https://huggingface.co/timdettmers/guanaco-7b. The proprietary models are GPT-3.5 Turbo (GPT-3.5-turbo-0125) [^14] https://openai.com/index/new-models-and-developer-products-announced-at-devday/, GPT-4 (GPT-4-0613), and Gemini-Pro [^15] https://ai.google.dev/.
Harmful questions collection. Following existing jailbreaking attacks [62], we employed a tailored subset of AdvBench Dataset. This subset encompasses a broad spectrum of prohibited scenarios, including illegal activities, immoral behavior, discriminatory content, and toxic material. This subset was carefully selected because it contain questions either manually crafted by the authors or generated through crowdsourcing, providing a close reflection of real-world scenarios.
Experimental metrics. We evaluate the effectiveness of Papillon through two metrics, i.e., Attack Success Rate (ASR) and Average Queries (AQ). To make a fair comparison, we calculate the AQ only for successful jailbroken questions. AQ can represent the query overhead of different attacks. Note that both Papillon and baseline attacks use GPT-3.5 turbo in prompt generation and jailbreak judgment procedures, but we use GPT-4 to evaluate attack performance in terms of the output’s relevance to the harmful query for a fairer and more accurate comparison. The length of the generated prompts is controlled using the $max\_new\_tokens$ hyperparameter. Specifically, for the “role-play" and “contextualization" mutations, this value is set to 200 tokens, while an additional 100 tokens are allocated for the “expand" mutation.
More details of the target LLMs and metrics are shown in the appendix.
Our experiments were conducted on a server equipped with 2 NVIDIA A100 PCIe GPUs, each with 40GB of memory. The server’s CPU is an Intel(R) Xeon(R) Silver 4210R with 10 cores, endowed with 188GB of memory. In terms of software, the server runs on the Ubuntu 22.04.1 LTS operating system. The experiments utilized Python version 3.10.13, CUDA version 12.1, PyTorch version 2.1.2, and the transformers library version 4.36.2.
5.2 Comparison with Baselines
We first compare Papillon with five state-of-the-art jailbreaking attacks, including AutoDAN [30], GCG [62], PAIR [7], Gptfuzzer [54], and TAP [34]. The comparison results are shown in Table 10.
We can see that Papillon can achieve the highest attack success rate compared to the baselines across all the victim LLMs. In terms of open-source models, Papillon can achieve ASRs of 100%, 58%, 100%, and 98% for Vicuna-7B-v1.3, Llama-2-7B-chat, Baichuan2-7B-chat, and Guanaco-7B, respectively, while the highest ASRs of the five baselines only achieve 90% (Gptfuzzer), 52% (TAP), and 58% (GCG), and 48% (TAP), respectively. For proprietary models, Papillon achieves an ASR of 90% for GPT-3.5 turbo, 80% for GPT-4, and 82% for Gemini-Pro, significantly outperforming the baselines. Notably, for GPT-4 and Gemini-Pro, Papillon surpasses the baselines by nearly 60%. Additionally, Papillon requires fewer than 25 query times in most cases to achieve the attack goal.
The lower performance of GCG is likely due to our budget being set at only 100, which is significantly less than the 500 used in the original paper. Additionally, the original paper’s judge method is relatively coarse and may result in false positives for successful jailbreaks. As a result, the use of a more stringent evaluation method in the experiments resulted in a lower success rate. Examples like AutoDAN for Llama-2-7B-chat have significantly lower average queries (AQ) and attack success rates (ASR) because they can only successfully jailbreak easier questions, even after many attempts. Note that AQ is calculated only for questions that were successfully jailbroken.
Since Papillon does not rely on any pre-existing jailbreaking templates to initiate the attack, these results demonstrate its effectiveness as a powerful jailbreaking method. The success of Papillon can be attributed to its fuzz testing framework, mutator, and judge module designs.
5.3 Ablation Studies
Impact of different mutators. In Papillon, we design three mutators: Role-play, Contextualization, and Expand. Unlike existing works that rely on manual effort to design them, we propose novel LLM-empowered mutation modules. To explore their effectiveness, we test the ASR and AQ when Papillon adopts different combinations of them, i.e., “Role-play", “Role-play and Contextualization", and all three mutators. The results are shown in Table 11 (appendix).
As the number of mutator types increases, Papillon achieves a higher ASR while still maintaining a low query number. For example, when targeting GPT-3.5 turbo, Papillon attains an ASR of 76% and an AQ of 17.66 when only using the Role-play mutator. However, it achieves an ASR of 86% and an AQ of 16 when the Scenario mutator is added. When using all the mutators, Papillon reaches an ASR of 90% and an AQ of 22.93.
Impact of pre-jailbreak. Since Papillon starts without existing jailbreaking templates, early queries are challenging due to limited seed availability. To address this, we introduce a pre-jailbreak phase with initial jailbreaking attempts (e.g., 5 times) for each problem. Papillon operates in two stages: pre-jailbreak and final-jailbreak. The impact of pre-jailbreak on Papillon is shown in Table 12 (appendix).
The results indicate that incorporating a pre-jailbreak process significantly boosts the attack success rate in most cases, with improvements of over 6% when targeting GPT-4.
Impact of different judges. We adopt a two-level judge module to determine whether a jailbreak prompt is successful. In this section, we explore the impact of each judge on the attack performance of Papillon. The results are shown in Table 13 (appendix).
The results indicate that the stricter the judge module, the more powerful the jailbreak prompts Papillon can generate to deceive the LLM into producing harmful content. Using only RoBERTa, which assesses whether the generated content is harmful but cannot verify its correlation to the question, may lead to ineffective jailbreak prompts when calculating the attack success rate. A jailbreak prompt will only be considered successful if it is both harmful and relevant to the question.
Impact of prompt length. Unlike existing works that overlook prompt length, we limit the prompt length to 200 tokens in Role-play and Contextualization and 100 tokens in Expand in Papillon to decrease attack costs and improve evasiveness. In this section, we explore the impact of prompt length on attack performance. We vary the prompt length from 50 to 300 tokens in Role-play and Contextualization, and the results are shown in Table 15 (appendix).
As the prompt length increases, the attack success rate improves, making the attack more effective. However, this also incurs higher costs since LLM API pricing is based on token length. This creates a trade-off between attack effectiveness and query cost. In Papillon, we set the default prompt length to 200 tokens. Even with a budget of 50 tokens, we achieve an ASR of 100% for Vicuna-7B-v1.3, 98% for Baichuan2-7B, 60% for GPT-3.5 turbo, and 46% for Gemini-Pro. Note that attackers can dynamically adjust the prompt length when using Papillon to launch the attack.
Table 2: Impact of the query budget of the pre-jailbreak.
Target model Metrics budget=5 budget=10 budget=15 budget=20
Vicuna-7B-v1.3 ASR ($\uparrow$) 100% 100% 100% 100%
AQ ($\downarrow$) 4.70 4.40 5.00 5.80
Llama-2-7B-chat ASR ($\uparrow$) 56% 58% 64% 56%
AQ ($\downarrow$) 40.86 37.41 42.75 31.75
Baichuan2-7B ASR ($\uparrow$) 100% 100% 100% 100%
AQ ($\downarrow$) 5.56 3.90 3.78 5.16
Guanaco-7B ASR ($\uparrow$) 96% 98% 96% 98%
AQ ($\downarrow$) 14.62 7.67 10.79 8.62
GPT-3.5 turbo ASR ($\uparrow$) 88% 90% 90% 94%
AQ($\downarrow$) 16.76 18.04 18.27 16.70
GPT-4 ASR ($\uparrow$) 74% 80% 86% 82%
AQ($\downarrow$) 37.08 27.20 23.09 32.90
Gemini-Pro ASR ($\uparrow$) 80% 82% 84% 96%
AQ($\downarrow$) 21.86 19.02 21.18 16.56
Table 3: Mutation with smaller models.
Target model Metrics Vicuna-7b-v1.3 Guanaco-7B
Vicuna-7B-v1.3 ASR ($\uparrow$) 100% 100%
AQ($\downarrow$) 2.86 6.69
Llama-2-7B-chat ASR ($\uparrow$) 58% 56%
AQ($\downarrow$) 36.97 34.68
Baichuan2-7B ASR ($\uparrow$) 100% 100%
AQ($\downarrow$) 3.16 6.30
Guanaco-7B ASR ($\uparrow$) 100% 100%
AQ($\downarrow$) 5.84 12.46
GPT-3.5 turbo ASR ($\uparrow$) 90% 92%
AQ($\downarrow$) 24.43 20.24
GPT-4 ASR ($\uparrow$) 80% 80%
AQ($\downarrow$) 31.28 29.47
Gemini-Pro ASR ($\uparrow$) 76% 74%
AQ($\downarrow$) 24.32 24.19
Table 4: Performance of Papillon when targeting larger open-source LLMs.
Target model ASR AQ
Vicuna-13B-v1.3 100% 12.36
Vicuna-33B-v1.3 100% 4.58
Llama-2-70B 100% 7.20
Llama-3.3-70B-Instruct 64% 30.81
Impact of different query budgets of pre-jailbreak. We investigate the effect of varying query budgets in the pre-jailbreak phase. Specifically, we examine query budgets ranging from 5 to 20 queries, with the results presented in Table 2. It is shown that increasing the query budget in the pre-jailbreak phase leads to a rise in the ASR of Papillon, but the improvement becomes less pronounced with higher budgets. For our experiments, we set the query budget for the pre-jailbreak phase to 10 by default before performing the final-jailbreak. Note that the attacker can choose the most suitable query budget according to the attack performance.
Impact of total query budget. We explore the impact of the query budget in Table 14 (appendix). While the attack success rate generally improves with a higher query budget due to the increased number of attempts to mislead the LLM, it does not continue to increase indefinitely. Some jailbreak prompts can still be circumvented by the LLM, even after more than 150 queries. Consistent with existing studies, we set the default query budget to 100. If prompts are ineffective after 100 queries, they are considered unsuccessful.
Mutation with smaller model. By default, we use GPT-3.5 Turbo for high-quality prompt generation. To evaluate the performance of Papillon with smaller models, we tested Vicuna-7B-v1.3 and Guanaco-7B, as shown in Table 3. While Papillon does not strictly rely on large LLMs, smaller models can lower costs, albeit with potential trade-offs in prompt quality and attack success rates. For instance, Papillon achieves a 100% ASR when targeting the proprietary Gemini-Pro model with GPT-3.5 Turbo, whereas using Guanaco-7B results in a reduced ASR of 74%.
Attack against larger open-sourced LLMs. Apart from the four open-sourced models, i.e., LLaMA-2-7B-chat, Vicuna-7B-v1.3, Baichuan2-7B, and Guanaco-7B, we also explore the effectiveness of Papillon against larger models, including Vicuna-13B-v1.3161616https://huggingface.co/lmsys/vicuna-13b-v1.3, Vicuna-33B-v1.3171717https://huggingface.co/lmsys/vicuna-33b-v1.3, and Llama-2-70B [^18] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf, and Llama-3.3-70B-Instruct [^19] https://build.nvidia.com/meta/llama-3_3-70b-instruct. The results are shown in Table 4. We can see that Papillon maintains a high attack success rate (even 100%) with a limited query budget, even when targeting larger LLMs, highlighting the robustness and power of Papillon.
5.4 Attack Transferability
We further investigate the transferability of Papillon. Transferability refers to the extent to which prompts designed to jailbreak the target model can successfully jailbreak another model.
LLM transferability. We first evaluate transferability by using jailbreak prompts generated by the source model to target another LLM. The results are shown in Table 17 (appendix).
It is evident that prompts generated by one model can sometimes successfully mislead other LLMs into generating malicious content. For example, jailbreak prompts generated by Baichuan2-7B achieve a 70% attack success rate against Vicuna-7B-chat. Overall, Papillon demonstrates notable attack transferability across different models.
MLLM transferability. We then assess the transferability of the PAPILLON-generated jailbreak prompts to multimodal large language models (MLLMs). Specifically, we evaluate two commonly used MLLMs, i.e., Llava-1.5-7B and Llava-1.5-13B. The models are provided with an image and a text prompt as inputs, generating text responses. In the experiments, successful jailbreaking templates are printed on white-background images and used as image inputs, while harmful questions serve as text prompt inputs to mislead the MLLMs into generating malicious or harmful responses. We test 50 jailbreak prompts, and the results are presented in Table 5. Notably, the successful prompts for the LLM are also effective in misleading the MLLMs in many instances, achieving an attack success rate of up to 76% for Llava-1.5-7b and 66% for Llava-1.5-13b.
6 Resiliency against Defenses
6.1 Perplexity Filter
Perplexity is typically low for benign prompts but tends to be high for jailbreak prompts, making it useful for detecting jailbreaking attacks. To reduce false negatives and false positives, Alon et al. [1] proposed a classifier based on two features: perplexity and token sequence length. Jailbreak prompts are typically lengthy, often including long suffixes or prefixes, whereas many regular prompts with high perplexity are notably short. As shown in Table 16 (appendix), when applying the perplexity filter [1] to Papillon, we observe that Papillon maintains a high attack success rate. The ASR drop of Papillon is less than 10% in nearly all cases. The potential reason is that the jailbreak prompts generated by Papillon maintain semantic coherence with low perplexity. This allows these jailbreak prompts to bypass the perplexity-based detection mechanism by closely mimicking regular, benign prompts, thus evading detection.
6.2 SmoothLLM
SmoothLLM [41] is based on the finding that adversarially generated prompts are fragile to character-level changes. It first introduces random input perturbations to multiple copies of a given input prompt and then aggregates the corresponding predictions to detect adversarial inputs. In our experiments, we employed the RandomSwapPerturbation method to ensure optimal defense performance. We set the perturbation percentage to 20% and utilized 10 copies of the swap method.
As shown in Table 16 (appendix), after integrating Smooth-LLM into Papillon, the high attack success rate is still high with only a slight drop. The potential reason is that Papillon is designed to be based on semantics rather than simple character matching. As a result, straightforward character replacement is ineffective at preventing Papillon.
Table 5: Transferability of the Papillon-generated jailbreak prompts to the multi-modal LLMs.
Source model Llava-1.5-7B Llava-1.5-13B
Vicuna-7B-v1.3 72% 66%
Vicuna-13B-v1.3 76% 64%
Llama-2-70B 76% 62%
6.3 Llama Guard
Llama Guard [17] is a safeguard model based on the Llama-2-7B architecture, designed to classify prompts and outputs in AI-human interactions as “safe" or “unsafe". Unsafe prompts are filtered to defend against jailbreak attempts. In our study, we targeted LLMs equipped with Llama Guard, with results presented in Table 6. The experimental findings reveal that we can achieve a high attack success rate even under Llama Guard’s defense. This is attributed to our ability to craft more complex and mutated prompts, which Llama Guard misclassifies as “safe," enabling us to bypass the safeguard.
6.4 Hybrid Defense
In addition to single defenses, we also consider two combinations of multiple defenses, referred to as hybrid defenses. More specifically, we consider two kinds of hybrid defenses, i.e., Perplexity filter + SmoothLLM and SmoothLLM + Llama Guard. In the former case, the defender initially employs a perplexity filter to screen prompts. Prompts that pass this filter are then subjected to the SmoothLLM defense. In the latter case, the defender first leverages Llama Guard to filter unsafe prompts and then applies the SmoothLLM defense to the prompts that pass the Llama Guard filter. If the output generated by the model after applying these hybrid defenses still meets the attack criteria, it is considered a successful jailbreak. The results are shown in Table 7.
Hybrid defenses generally provide stronger protection than single defenses. However, Papillon still achieves a high attack success rate in such cases. This success is due to Papillon’s reliance on its semantic prompts with low perplexity, rather than simplistic character matching. These semantic prompts are not only complex but also undergo deliberate mutations, allowing Papillon to generate highly dynamic and contextually adaptive inputs that effectively bypass detection mechanisms.
Table 6: Apply Llama Guard defense to Papillon.
Target model Papillon After Llama Guard
Vicuna-7B-v1.3 100% 96%
Llama-2-7B-chat 58% 58%
Baichuan2-7B 100% 98%
Guanaco-7B 98% 96%
GPT-3.5 turbo 90% 64%
GPT-4 80% 64%
Table 7: Apply hybird defense to Papillon.
Target model Papillon Perplexity + SmoothLLM SmoothLLM + Llama Guard
Vicuna-7B-v1.3 100% 86% 100%
Llama-2-7B-chat 58% 56% 54%
Baichuan2-7B 100% 92% 98%
Guanaco-7B 98% 86% 84%
GPT-3.5 turbo 90% 80% 58%
GPT-4 80% 80% 52%
7 Human Evaluation
Threshold hyperparameter. To validate the selection of the threshold hyperparameter in the LLM-based judge module, we conducted a human evaluation on a subset of the AdvBench Dataset. We generate jailbreak prompts with different scores under the Papillon framework. A total of 30 volunteers participated in this evaluation by providing binary feedback on whether a given prompt is a successful jailbreaking instance or not. The results are summarized in Table 8. We observed that for a score of 7, a generated prompt was still misaligned with human evaluation. For a score of 8 or higher, all outputs were consistently aligned with human evaluation, achieving a 100% alignment rate. This indicates that a score of 8 serves as an appropriate threshold for distinguishing jailbroken outputs that meet human evaluation standards. Based on these findings, we adopt a threshold value of 8 for our experiments.
Prompts quality. To evaluate the interpretability and naturalness of the generated prompts, we developed two criteria for quantifiable human evaluation. For interpretability, we measure how easily humans can understand the mechanism behind the jailbreak prompt. In simple terms, it assesses whether the evaluator can easily comprehend why the prompt triggers the jailbreak effect. Naturalness is assessed based on the readability of the prompt. Both criteria are scored on a scale of 1 to 5, with higher scores indicating better performance. A total of 30 volunteers participated in this evaluation. For the evaluation, we selected 50 manually crafted jailbreak prompts from Gptfuzzer [54] and 50 jailbreak prompts generated by Papillon, corresponding to the same set of questions. For fairness, we ensured that the formats of the artificial and generated prompts were consistent. Additionally, we randomized the order of the 50 manual prompts and our 50 generated prompts to avoid regularity that could influence the evaluation process. The results are summarized in Table 9. We observed that our generated jailbreak prompts match or surpass manually crafted ones in interpretability and naturalness. Papillon can achieve a high ASR while maintaining outstanding prompt quality.
Table 8: Results of human evaluation to determine the optimal threshold hyperparameter on the AdvBench subset.
Score 6 7 8 9 10
# of generated prompts (Papillon) 0 3 15 12 139
# of judged as jailbroken prompts (Human) 0 2 15 12 139
Alignment rate - 66% 100% 100% 100%
Table 9: Human evaluation to the quality of manual prompts and Papillon generated prompts.
Quality Interpretability Naturalness
# of judged as jailbroken prompts (Human) 3.98 3.66
# of generated prompts (Papillon) 4.64 4.60
All Prompts 4.31 4.03
8 Conclusion and Future Work
This paper introduces Papillon, a fuzz testing-driven framework designed to jailbreak large language models. Papillon can automatically generate jailbreak prompts with only black-box access to the target LLMs. Unlike existing methods that rely on predefined jailbreaking templates, Papillon begins with an empty seed pool, eliminating the need for any related jailbreaks. To enhance attack performance, we design three novel question-dependent mutation strategies that generate semantically meaningful and concise mutations. We also proposed to use a two-level judge module to distinguish successful jailbreaks. Extensive experiments on 7 representative LLMs demonstrate the effectiveness of Papillon compared to 5 state-of-the-art attacks. Papillon also shows resistance to advanced defenses and exhibits high transferability.
There are various potential areas for future exploration. First, while we designed three effective mutators in Papillon, there is still room to develop additional mutators for the mutator pool, such as rephrasing or using ciphertext. These potential mutators may further improve the attack success rate. Second, we will evaluate other state-of-the-art prompt compression methods, such as [36] and [59], to determine if they can generate more effective jailbreak prompts compared to our LLM-driven prompt shortening techniques. Third, effective defense against Papillon is necessary to reduce the potential risks of such attacks. Fourth, Papillon mainly focuses on jailbreaking attacks using English words and targets only English models. An interesting question is whether Papillon can also be effective for other language LLMs. We will explore this in the future. Finally, our findings reveal a critical limitation in aligned LLMs: they are effective in generating jailbreaking templates that do not necessarily exhibit explicit malicious intent. This raises significant safety concerns, indicating the need for further research on alignment strategies to mitigate such behaviors.
Acknowledgement
This research is supported by the National Research Foundation, Singapore under its Strategic Capability Research Centres Funding Initiative and Infocomm Media Development Authority under its Future Communications Research & Development Programme. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore. Yanjiao Chen’s work was partially supported by Ant Group through CCF-Ant Research Fund. Qian Wang’s work was partially supported by the NSFC under Grants U2441240 (“Ye Qisun” Science Foundation), 62441238 and U21B2018. The first two authors contributed equally to this work. Chen Chen is the corresponding author.
Ethics Considerations
In developing the Papillon framework, we have carefully considered the ethical implications of our research. This research inherently involves the potential generation of prompts that could lead to harmful or offensive content. However, we have adopted several measures to ensure that our findings are handled ethically and responsibly.
Stakeholders and Potential Impact. The key stakeholders involved in our research include LLM developers, users of LLM-powered systems, and the wider public who might be indirectly impacted by misuse of these technologies. The release of the Papillon framework is intended to assist researchers and developers in identifying and addressing vulnerabilities in LLMs, thereby improving their overall security and contributing to safer and more robust AI systems. However, we acknowledge the risk that malicious actors could use Papillon to bypass existing defenses, potentially leading to harmful applications. This risk is further amplified by the possibility that Papillon could be adapted to attacking multimodal LLMs, potentially inducing outputs that violate their intended ethical constraints, such as generating images with harmful content (e.g., CSAM, non-consenual sexual material, etc.). To mitigate these risks, we emphasize the need for responsible use of the framework, explicitly discourage any unethical applications, and provide clear guidelines to minimize the likelihood of misuse.
Responsible Disclosure and Dual-Use Concerns. Generating harmful outputs presents significant ethical challenges. To address this, we setup a restricted environment to generate harmful prompts during experiments, only allowing the research team to access. While the core components of Papillon will be open-sourced, specific prompts used for successful jailbreaks will not be released to prevent misuse. This approach balances transparency and research utility with minimizing risks of abuse. Additionally, we will actively monitor community feedback and adapt dissemination practices as needed to prevent potential misuse.
Protection of Research Team Members. The research team has been mindful of the psychological and ethical implications of working with potentially disturbing content. We have ensured that all team members are aware of the risks and have access to support if needed. We have established protocols for reporting any distress experienced by team members and provided access to mental health resources.
By adhering to these principles, we aim to ensure that our research on jailbreaking attacks remains within ethical boundaries and contributes to the responsible advancement of AI technology. We hope that our work not only highlights vulnerabilities but also serves as a catalyst for developing stronger defenses, ultimately contributing to a safer deployment of LLMs.
Compliance with the Open Science Policy
We are committed to the Open Science Policy and will make our research artifacts, such as code and non-sensitive data, publicly available upon publication. This ensures that our work can be replicated and verified by others in the community. Specifically, the following artifacts will be made publicly available upon acceptance: 1) The full implementation of the Papillon framework, including all components necessary for replication, such as the fuzz-testing module, mutation strategies, and the two-level judge module; 2) Detailed scripts and configuration files for running the experiments and reproducing all results presented in the paper. This includes scripts for evaluating attack success rates, processing intermediate results, and running comparative baselines.
To prevent misuse, we will not publicly release any harmful content, such as specific jailbreak prompts. These prompts, as well as any outputs containing potentially harmful responses, will remain restricted to the research team and will only be shared with verified researchers upon request. Access requests will be subjected to a strict review process to ensure ethical use. Our goal is to balance transparency with responsibility, safeguarding against potential harm while enabling meaningful scientific progress.
Our submission includes a clearly-marked section on ethical considerations, and we have followed the conference’s ethical guidelines throughout our research. By doing so, we aim to contribute to open science while safeguarding against potential harms.
Appendix A Victim Large Language Models
Vicuna-7b-v1.3. Vicuna-7b-v1.3 is developed by LMSYS through the process of fine-tuning the LLaMA model using user-shared conversations from ShareGPT. Vicuna-7b is equipped with approximately 7 billion parameters.
LLaMA-2-7b-chat. The LLaMA-2 series represents a family of LLMs developed by Facebook. It is open-sourced, and we focus on the LLaMA-2 7b model, which comprises approximately 7 billion parameters.
Baichuan2-7B-Chat. Baichuan2-7B is a large-scale, open-source language model developed by Baichuan Intelligence Inc. With 7 billion parameters, Baichuan2-7B is well-suited for applications like text generation, translation, summarization, and question answering.
Guanaco-7B. The Guanaco models are open-source, fine-tuned chatbots created by applying 4-bit QLoRA tuning to LLaMA base models using the OASST1 dataset. Guanaco-7B is equipped with approximately 7 billion parameters.
GPT-3.5 turbo and GPT-4. GPT-3.5 turbo and GPT-4 are large language models developed by OpenAI. They are accessible exclusively through a black-box API.
Gemini-Pro. The Gemini series is provided by Google. Gemini-Pro is their most advanced model, and it has an API available to the public. It is also only accessible through a black-box API.
Table 10: Comparison of Papillon with GCG [62], PAIR [7], AutoDAN [30], Gptfuzzer [54], and TAP [34]. Since GCG and AutoDAN are white-box jailbreaking attacks, they cannot be evaluated on proprietary LLMs such as GPT-3.5, GPT-4, and Gemini-Pro.
Target model Metrics AutoDAN GCG PAIR Gptfuzzer TAP Papillon
Vicuna-7B-v1.3 ASR ($\uparrow$) 16% 10% 72% 90% 16% 100%
AQ($\downarrow$) 22.88 100 10.87 6.53 9.25 4.40
Llama-2-7B-chat ASR ($\uparrow$) 4% 2% 18% 38% 52% 58%
AQ($\downarrow$) 1.00 100 31.54 22.92 29.50 37.41
Baichuan2-7B ASR ($\uparrow$) 36% 58% 32% 40% 30% 100%
AQ ($\downarrow$) 58.67 100 39.14 47.81 43.53 3.90
Guanaco-7B ASR ($\uparrow$) 20% 14% 24% 18% 48% 98%
AQ($\downarrow$) 6.80 100 29.45 23.65 51.33 7.67
GPT-3.5 turbo ASR ($\uparrow$) - - 52% 88% 6% 90%
AQ($\downarrow$) - - 23.70 19.98 20.67 18.04
GPT-4 ASR ($\uparrow$) - - 12% 22% 2% 80%
AQ($\downarrow$) - - 16.67 19.63 6.00 27.20
Gemini-Pro ASR ($\uparrow$) - - 12% 10% 4% 82%
AQ($\downarrow$) - - 21.89 33.16 77.00 19.02
  • ($\uparrow$) signifies that a higher value is preferable, while ($\downarrow$) indicates that a lower value is more desirable.
[…]
Appendix B Experimental Metrics
Attack success rate (ASR). As talked about in Section 4.2, we design a two-level judge module to judge the results of various jailbreaking attacks. Specifically, a jailbreak is considered successful if it passes the first judge model and scores greater than 8 in the second judge model. The attack success rate (ASR) is the ratio of successfully judged jailbreaks to the total number of questions.
$ASR=\frac{N_{S}}{N_{T}}\times 100\%,$ (7)
where $N_{S}$ is the number of successful jailbreaks, and $N_{T}$ is the total number of questions.
Average queries (AQ). To make a fair comparison, we calculate the average Queries (AQ) only for successful jailbroken questions. Average Queries (AQ) can represent the query overhead of different attacks.
$\text{AQ}=\frac{\sum_{i=1}^{N_{s}}Q_{i}}{N_{s}}$ (8)
where $Q_{i}$ is the number of queries for the $i$-th successfully jailbroken question, and $N_{s}$ is the total number of successfully jailbroken questions.
Appendix C Explanation of Some Terminologies
Round Robin. Round Robin cycles through the seed pool, ensuring comprehensive exploration.
UCB. UCB assigns a score to each seed, with the highest-scoring seed being selected.
MCTS-Explore. MCTS-Explore is a variant of the Monte Carlo Tree Search (MCTS) algorithm that balances the efficiency and diversity of seed selection.
Generate. Generate creates template variations that retain a consistent style while incorporating different content.
Crossover. Crossover merges two different jailbreaking templates into a novel template.
Expand. Expand expands the content of an existing template by adding supplementary material.
Shorten. Shorten condenses a template to ensure it remains meaningful yet more succinct.
Rephrase. Rephrase restructures the given template to maximize semantic preservation while altering its phrasing.
Table 11: Impact of using different mutators.
Target model Metrics Role-play Role-play+Contextualization All
Vicuna-7B-v1.3 ASR ($\uparrow$) 100% 96% 100%
AQ($\downarrow$) 7.20 6.14 4.40
Llama-2-7B-chat ASR ($\uparrow$) 24% 50% 58%
AQ($\downarrow$) 43.00 32.04 37.41
Baichuan2-7B ASR ($\uparrow$) 84% 90% 100%
AQ($\downarrow$) 4.72 2.29 3.90
Guanaco-7B ASR ($\uparrow$) 88% 96% 98%
AQ($\downarrow$) 14.64 9.69 10.82
GPT-3.5 turbo ASR ($\uparrow$) 76% 86% 90%
AQ($\downarrow$) 17.66 16.00 22.93
GPT-4 ASR ($\uparrow$) 34% 46% 80%
AQ($\downarrow$) 49.41 46.87 27.20
Gemini-Pro ASR ($\uparrow$) 54% 62% 82%
AQ($\downarrow$) 32.30 28.74 25.03
  • ($\uparrow$) signifies that a higher value is preferable, while ($\downarrow$) indicates that a lower value is more desirable.
Table 12: Impact of pre-jailbreak process.
Target model Metrics No pre-jailbreak Pre-jailbreak
Vicuna-7B-v1.3 ASR ($\uparrow$) 100% 100%
AQ ($\downarrow$) 4.46 4.40
Llama-2-7B-chat ASR ($\uparrow$) 46% 58%
AQ ($\downarrow$) 23.57 37.41
Baichuan2-7B ASR ($\uparrow$) 100% 100%
AQ($\downarrow$) 2.48 3.90
Guanaco-7B ASR ($\uparrow$) 98% 98%
AQ ($\downarrow$) 9.86 7.67
GPT-3.5 turbo ASR ($\uparrow$) 86% 90%
AQ ($\downarrow$) 21.65 22.93
GPT-4 ASR ($\uparrow$) 74% 80%
AQ($\downarrow$) 19.78 27.20
Gemini-Pro ASR ($\uparrow$) 72% 82%
AQ($\downarrow$) 18.86 25.03
Table 13: Impact of using different judges.
Target model Metrics RoBERTa RoBERTa+ChatGPT
Vicuna-7B-v1.3 ASR ($\uparrow$) 94% 100%
AQ ($\downarrow$) 2.09 4.40
Llama-2-7B-chat ASR ($\uparrow$) 36% 58%
AQ ($\downarrow$) 31.00 37.41
Baichuan2-7B ASR ($\uparrow$) 90% 100%
AQ($\downarrow$) 1.80 3.90
Guanaco-7B ASR ($\uparrow$) 88% 98%
AQ ($\downarrow$) 5.36 10.82
GPT-3.5 turbo ASR ($\uparrow$) 74% 90%
AQ ($\downarrow$) 14.95 22.93
GPT-4 ASR ($\uparrow$) 40% 80%
AQ($\downarrow$) 26.90 27.20
Gemini-Pro ASR ($\uparrow$) 42% 82%
AQ($\downarrow$) 25.57 25.03
Table 14: Impact of the query budget on the attack performance.
Target model Metrics budget=50 budget=75 budget=100 budget=125 budget=150
Vicuna-7B-v1.3 ASR ($\uparrow$) 100% 98% 100% 100% 100%
AQ($\downarrow$) 5.00 3.52 5.76 5.96 5.42
Llama-2-7B-chat ASR ($\uparrow$) 38% 38% 58% 62% 64%
AQ($\downarrow$) 16.68 19.79 37.41 36.08 18.06
Baichuan2-7B ASR ($\uparrow$) 100% 100% 100% 100% 100%
AQ ($\downarrow$) 3.48 3.79 4.42 5.47 4.10
Guanaco-7B ASR ($\uparrow$) 94% 96% 98% 98% 98%
AQ($\downarrow$) 9.06 8.60 7.67 14.88 7.51
GPT-3.5 turbo ASR ($\uparrow$) 74% 80% 90% 90% 94%
AQ($\downarrow$) 10.81 15.51 22.93 26.60 29.15
GPT-4 ASR ($\uparrow$) 54% 56% 80% 76% 78%
AQ($\downarrow$) 15.85 23.70 27.20 35.99 39.97
Gemini-Pro ASR ($\uparrow$) 60% 56% 82% 72% 68%
AQ($\downarrow$) 14.17 17.36 25.03 24.88 26.56
  • ($\uparrow$) signifies that a higher value is preferable, while ($\downarrow$) indicates that a lower value is more desirable.
Table 15: Impact of the prompt length on the attack performance.
Target model Metrics 50 tokens 100 tokens 150 tokens 200 tokens 250 tokens 300 tokens
Vicuna-7B-v1.3 ASR ($\uparrow$) 100% 100% 100% 100% 100% 100%
AQ($\downarrow$) 5.86 5.50 3.58 4.40 6.10 4.50
Llama-2-7B-chat ASR ($\uparrow$) 50% 40% 44% 58% 60% 50%
AQ($\downarrow$) 32.00 20.05 12.45 37.41 34.17 24.00
Baichuan2-7B ASR ($\uparrow$) 98% 100% 100% 100% 100% 100%
AQ ($\downarrow$) 5.55 5.54 5.34 3.90 4.32 4.54
Guanaco-7B ASR ($\uparrow$) 100% 100% 100% 98% 98% 98%
AQ($\downarrow$) 5.98 6.90 9.48 7.67 10.55 12.14
GPT-3.5 turbo ASR ($\uparrow$) 60% 70% 82% 90% 94% 88%
AQ($\downarrow$) 28.53 19.11 16.88 18.04 16.15 21.00
GPT-4 ASR ($\uparrow$) 36% 78% 66% 80% 72% 76%
AQ($\downarrow$) 43.17 28.72 26.30 27.20 35.89 38.81
Gemini-Pro ASR ($\uparrow$) 46% 68% 74% 82% 76% 52%
AQ($\downarrow$) 20.87 24.26 22.89 25.03 32.16 17.27
  • ($\uparrow$) signifies that a higher value is preferable, while ($\downarrow$) indicates that a lower value is more desirable.
Table 16: Apply perplexity filter and SmoothLLm defense to Papillon.
Target model Papillon After perplexity filter After SmoothLLm
Vicuna-7B-v1.3 100% 96% 100%
Llama-2-7B-chat 58% 58% 56%
Baichuan2-7B 100% 96% 100%
Guanaco-7B 98% 88% 98%
GPT-3.5 turbo 90% 80% 70%
GPT-4 80% 80% 78%
Gemini-Pro 82% 72% 58%
Table 17: Cross-model transferability.
Source model Vicuna-7B-v1.3 Llama-2-7B-chat Baichuan2-7B Guanaco-7B GPT-3.5 turbo GPT-4 Gemini-Pro
Vicuna-7B-v1.3 100% 10% 70% 24% 22% 6% 12%
Llama-2-7B-chat 46% 58% 40% 34% 22% 14% 28%
Baichuan2-7B 70% 4% 100% 34% 22% 8% 10%
Guanaco-7B 68% 12% 68% 98% 26% 16% 22%
GPT-3.5 turbo 62% 12% 56% 34% 90% 10% 22%
GPT-4 76% 2% 78% 46% 66% 80% 34%
Gemini-Pro 40% 12% 42% 30% 28% 16% 82%